OpenStudy (lacris):

2x-3[2x-(3-4x)] = ?? I need to simplify this but I keep getting 12x+9

OpenStudy (agent0smith):

2x-3[2x-(3-4x)] = 2x-3[2x-3+4x)] = 2x-3[6x-3] = 2x-18x+9

OpenStudy (lacris):

oh my gosh I see what I did now! I multiplied 6x and -3 to each number in the brackets, thank you so much for helping me!

OpenStudy (lacris):

you just had to multiply the -3 to the bracket? Not the 2x as well?

OpenStudy (lacris):



OpenStudy (lacris):

ahh, I see there are no brackets around the 2x-3, my bad, thanks for your patience :)